Find the value of (x+Y)if x+Y=5 and x-Y=15


1
Expert's answer
2022-01-31T13:14:59-0500

"x+y=5, x-y=15"

Then


"2x=20=>x=10"

"y=5-10=-5"

"(x, y)=(10, -5)"


"x+y=5"